OpenStudy (anonymous):

Evaluate each expression. log9(27)

OpenStudy (anonymous):

is it log base 9

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

you can use the property: \[\log_{b^{n}}(x) = \frac{1}{n}\log_b(x)\] this changes your problem to: \[\log_9(27) = \log_{3^{2}}(27)=\frac{1}{2}\log_3(3^{3}) = \frac{1}{2}*3 = \frac{3}{2}\]

OpenStudy (anonymous):

Simply u can write log9(base 27) = ln(3^3)/ln(3^2) den powers come out nd we get 3/2 ln3/ln3,,,, which is equal to 3/2